What charge is stored in a 180.0-"\mu"F capacitor when 120.0 V is applied to it







1
Expert's answer
2022-03-06T15:18:15-0500

The charge is given as the product of capacitance and voltage:


"Q= CV = 180\times 10^{-6}F\cdot 120V = 2.160\times 10^{-2}C"

Answer. "2.160\times 10^{-2}C".